Circular No. 7891

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cs IN NGC 6702 Independent reports have been received of the discoveries of an appar-

ent supernova on unfiltered CCD images obtained by LOTOSS (cf. IAUC 7514; via M. Ganeshalingam and W. D. Li, University of California at Berkeley) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) and by M. Armstrong, Rolvenden, England, with a 0.36-m reflector. SN 2002cs is located at α = 18^{h}46^{m}59^{s}.96, δ = +45^{o}42′19′′.7 (equinox 2000.0), which is 24′′.8 east and 0′′.8 south of the nucleus of NGC 6702 (measurements by Li; Armstrong gives the offsets as 24′′ east, 1′′ north). Unfiltered CCD magnitudes: 2001 Oct. 3.2 UT, [19.0: (KAIT); 2002 May 5.133, 16.3 (Armstrong); 5.5, ∼ 15.9 (KAIT); 6.3, ∼ 15.6 (KAIT). Armstrong adds that there is no trace of SN 2002cs on an earlier image taken by himself or on second Palomar Sky Survey blue, red, and infrared images.

SUPERNOVA 2002cr IN NGC 5468 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cr (cf. IAUC 7890), obtained by P. Berlind on May 3.27 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova before maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 2845 km/s for the host galaxy, the supernova expansion velocity is ∼ 12400 km/s for Si II (rest 635.5 nm) and 15600 km/s for Ca II (rest 395.1 nm). The spectral-feature age of the supernova (cf. IAUC 7774) is 5 ± 2 days before maximum light.

P. Meikle, Imperial College, London, reports that K.-H. Mack, and M. Marcha obtained a spectrum (range 382–722 nm, resolution 0.7 nm) of SN 2002cr on May 4.02 UT with the Isaac Newton Telescope at La Palma. Comparison of the spectrum with those of SNe 1994D and 1998bu indicates that SN 2002cr is a typical type-Ia event, ∼ 5 days before maximum light. The maximum expansion velocity measured from the blue wing of the Si II absorption is 19100 km/s (after correcting for the redshift of the host galaxy). A second spectrum obtained on May 5.06 by the same observers shows relatively little change in the spectral features.

Unfiltered CCD magnitudes: Apr. 16.4 UT, [19.0: (W. D. Li, University of California at Berkeley, 0.8-m Katzman Automatic Imaging Telescope); May 2.38, ∼ 15.9 (Li; independent discovery by LOTOSS); 3.716, 15.6 (L. A. G. Monard, Pretoria, S. Africa); 4.913, 15.2 (Monard); 5.639, 15.8 (R. Kushida, Yatsugatake South Base Observatory); 5.764, 15.0 (Monard).

2002 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green